# secReferences in the Online Appendix

### v1

@@ 4.1.1 Shareholders
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately affect a firm’s behavior, without having to wait a year to advocate for change at the next AGM.
@@ 4.1.2 Consumers
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value rebases a firm’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). That is, we investigate potential consumer boycotts in the months following an NGO campaign. Standard errors are clustered at the firm level.
@@
NGO campaigns on an AGM date impact both shareholders, who are more likely to make related proposals in future years, and consumers, who are more likely to boycott the firm in the following months.
@@ 5 Timing Campaigns to Influence AGMs
$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM. The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year and topic-by-year fixed effects, thereby exploiting only variation across campaigns and topics within firms. The standard errors are clustered by firm.
@@
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (e.g., Shu, 2024). Those recommendations have a critical influence on the vote decisions. We thus include management and ISS’s recommendations as covariates in (4).
@@
Note: E&S denote environmental and social proposals. The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0. The regressions also include firm and year-by-industry fixed effects. Standard errors are clustered at the firm level.
@@ 6.1 Campaign Timing and Accumulated NGO Visibility
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise. Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation, we include NGO-by-year fixed effects. The regression also includes firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. The standard errors are clustered at the NGO level. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample.
To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^8] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year.
@@
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time, thereby excluding selection on observables. For instance, Appendix Figures B7 and B8 show that NGO experience does not influence the choice of targets in terms of size (market capitalization), and Environmental and Social scores for Environmental and Social campaigns, respectively.
@@ 6.2 NGO Success and Reputation
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+u_{i,n,y},$
where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the firm-by-year, ngo-by-year, and topic-by-year through fixed effects. Standard errors are clustered by firm.
@@
Campaign on Same Topic 0.018 0.083∗∗ 0.046 0.076
@@
Campaign on Same Topic $\times$ Experience 0.024
AGM Experience (0/1) -0.036 -0.035
@@
Campaign on Same Topic $\times$ AGM Experience (0/1) 0.141∗∗ 0.126
@@
Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.054
AGM Experience -0.019
@@
Campaign on Same Topic $\times$ AGM Experience 0.068∗∗
@@
Column (1) of Table 1 reports the estimates for $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a topic being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
@@ 7 Discussion
Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions. In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaign against a target firm’s practice or strategy (Result 1). By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2). The goal of these campaigns is to alter a firm’s strategy by damaging its reputation (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
@@
Therefore, NGOs face a trade-off, which they navigate by adapting strategies over time. NGOs with limited activism history tend to campaign on the AGM date. These NGOs may favor early campaigns to build a reputation, as without it, they may be unable to influence shareholder votes. In contrast, more established NGOs prefer targeting firms just before their AGM to directly affect shareholder votes (Result 4), consistent with reputation as a signaling device.
@@
The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results. Thus, there are diminishing returns to reputation from consistently targeting AGMs. We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass. However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns, decreases support for shareholder proposals (Result 5).
@@
In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019). However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited because of limited information on NGOs. This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals. Similar approaches have been successfully used to identify the preferences of other stakeholders and may help reframe the role of AGMs not just as voting venues but also as forums where firms engage with and negotiate alongside outside stakeholders in ways that are not often observed by researchers but can be inferred from the underlying system of incentives (Fioretti et al., 2023). We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to that employed in this paper to quantitatively study the incentives for collaboration.
@@
By focusing on the timing of NGO campaigns, we demonstrate how this seemingly innocuous decision has significant implications for both targeted firms and the success of future initiatives. We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses. Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize their externalities.
@@ 8 Conclusion
This paper examines the timing and motives behind NGO campaigns, revealing that these campaigns are not randomly timed but often strategically target firms’ Annual General Meetings. While this timing does not affect votes on environmental and social proposals, which are cast beforehand, it enhances the NGO’s visibility, leading to greater consumer reactions and increasing the likelihood of related shareholder proposals in the following year. Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals. Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its lifecycle, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
@@ Appendix A Omitted Tables
Table A2: Campaign timing and target selection
@@ Appendix B Omitted Figures
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
@@
Note: This figure plots the estimated coefficients from the regressions of on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Standard errors are clustered at the firm level. The number of observations is 1,338.
@@
Note: This figure plots the estimated coefficients from the regressions of on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.

### v2

@@ 4.1.1 Shareholders
Importantly, we can rule out the possibility that NGOs partnering with shareholders drive this outcome. If that were the case, submitting a shareholder proposal followed by an early campaign would allow the partnering shareholders and NGOs to immediately gain broad support on the shareholder’s proposal, without having to wait a year to advocate for change at the next AGM.
@@ 4.1.2 Consumers
where $q$ indicates a quarter and $\alpha_{n}$ and $\alpha_{s(n),q}$ are fixed effects at either the firm level or the industry-by-quarter-by-year level, respectively. The dependent value rebases $n$’s sales by its asset to account for time variation in relative firm sizes, which is large in our sample (10 years, largest 3,000 firms). That is, we investigate potential consumer boycotts in the months following an NGO campaign. Standard errors are clustered at the firm level.
@@
NGO campaigns on an AGM date impact stakeholders. (i) Shareholders are more likely to make related proposals in future years. (ii) Consumers are more likely to boycott the firm in the following months.
@@ 5 Timing Campaigns to Influence AGMs
$\text{Success}_{i,n,m,y}=\sum^{0}_{\begin{subarray}{c}l=-6\end{subarray}}\beta_{l}\cdot\text{Campaign}_{i,n,m+l,y}+\alpha_{n,y}+\alpha_{i,y}+\alpha_{m}+u_{i,n,m,y},$ (4)
where $\text{Success}_{i,n,m,y}$ is a dummy indicating the success of proposal $i$ at firm $n$ in month $m$ in year $t$, and $\text{Campaign}_{i,n,m+l,t}$ is a dummy indicating the occurrence of an NGO campaign on the same topic as proposal $i$, $l$ months before the AGM. The reference period covers the months that occur more than six months before the AGM of year $y$—i.e., $m\in[-11,-7]$. We account for firm-by-year, and topic-by-year fixed effects, as well as seasonality with month fixed effects, thereby exploiting only variation across campaigns and topics within firm-year pairs. The standard errors are clustered by firm.
@@
To understand why campaigns 6 months before the AGM are also somewhat successful, we estimate (4) on the subset of voted proposals, with the percentage of votes in favor as the dependent variable. Note that every proposal to be voted on at the AGM is accompanied by a vote recommendation by the management (“For”, “Against”, or “Abstain”). Institutional investors can also follow the recommendations of proxy advisors, the most popular one being ISS (Shu, 2024). Those recommendations have a critical influence on the vote decisions. We thus include management and ISS’s recommendations as covariates in (4).
@@
Note: E&S denote environmental and social proposals. The figure plots the estimated coefficients of regressions of a dummy indicating that a shareholder proposal is withdrawn or passes if it goes to a vote (left panel, based on 911 proposals), or the percentage of votes in favor of the proposal when it goes to the vote (right panel, based on 2,078 proposals), on dummies indicating the occurrence of an NGO campaign on the same topic $k$ months before the AGM, for $k$ varying from -6 to 0. The regressions also include month, firm-by-year, and topic-by-year fixed effects. Standard errors are clustered at the firm level.
@@ 6.1 Campaign Timing and Accumulated NGO Visibility
$\text{Distance to AGM}_{c(i,n,g),y}$ $=\sum_{\begin{subarray}{c}\tau>1\end{subarray}}\beta_{\tau}\cdot\mathbf{1}_{\{\text{Campaign \#}_{c(\cdot),g,y}=\tau\}}+\alpha_{g,y}+\alpha_{n,y}+\alpha_{i,y}+u_{c(g,n,i),y},$ (5)
which examines $\text{Distance to AGM}_{c(i,n,g),y}$, the distance (in months) of campaign $i$ to the AGM of targeted firm $n$, as a function of the number of past campaigns by NGO $g$, using a dummy, $\text{Campaign \#}_{c,g,y}$, that is one if the number of NGO campaigns launched by NGO $g$ until year $y$ is equal to $\tau$ and zero otherwise.
Since our campaign data start in 2010, we do not know how many campaigns an NGO launched before 2010. To control for this truncation as well as time varying confounders—e.g., the size of the NGO social media network across years (Tufekci, 2017)—we include NGO-by-year fixed effects. We also include firm-by-year and topic-by-year fixed effects to account for firm and topic-specific occurrences that, varying over time, may incentivize an NGO to launch a campaign. Hence, the identification of $\beta_{\tau}$ comes from multiple campaigns $c(\cdot)$ by the same NGO $g$ within year $y$ [^9] In Appendix Figure B6, we present similar results when controlling only for firm and NGO fixed effects, so that the estimation of $\hat{\beta}_{\tau}$ come from variation across both campaigns and years. Because of the scope of this analysis, we only consider NGOs with at least two campaigns, but the results still hold in the full sample. The standard errors are clustered at the NGO level.
To interpret the estimated $\beta_{\tau}$ coefficients, consider that, on average, the first NGO campaign is exactly on the month of the AGM. Thus, Figure 4 indicates that as the NGO’s experience increases, the new campaigns are launched before the AGM date. [^10] The large negative coefficients estimated for NGOs with many campaigns—e.g., more than 9—are driven by outlier NGOs with several campaigns per year. We postpone the interpretation of this result to Section 7, which sums up and discuss the results.
@@
Alternative mechanism—selection. NGO do not increasingly launch campaigns against firms that may be more easily influenced over time. Appendix Figures B8 and B9 exclude selection on observables. For instance, target sizes do not covary (market capitalization) with NGO experience. In addition, environmental (social) scores do not covary with an NGO’s experience in environmental (social) campaigns.
@@ 6.2 NGO Success and Reputation
$+\alpha_{n,y}+\alpha_{g,y}+\alpha_{i,y}+\alpha_{m(c(i,n,g))}+u_{i,n,y},$
where $\text{Success}_{i,n,y}$ is defined as in (4)—namely, either a successful vote or a withdrawn proposal. On the right-hand side, $\text{Campaign on Same Topic}_{i,n,y}$ is one if there has been a campaign by NGO $g$ against firm $n$ on topic $i$ in the ten months preceding the AGM and zero otherwise, and $\text{Experience}_{g,y}$ indicates the number of campaigns NGO $g$ had launched in the past. We control for all confounding variations at the month, firm-by-year, NGO-by-year, and topic-by-year through fixed effects. Standard errors are clustered by firm.
@@
Campaign on Same Topic 0.008 0.076 0.032 0.070
@@
Campaign on Same Topic $\times$ Experience 0.026
AGM Experience (0/1) -0.042 -0.041
@@
Campaign on Same Topic $\times$ AGM Experience (0/1) 0.169∗∗ 0.150∗∗
@@
Campaign on Same Topic $\times$ Non-AGM Experience (0/1) 0.065
AGM Experience -0.018
@@
Campaign on Same Topic $\times$ AGM Experience 0.074∗∗
@@
Column (1) of Table 1 reports the estimated $\{\beta_{1},\beta_{2},\beta_{3}\}$. Campaigning on the same topic as a shareholder proposal only weakly influences the outcome of the proposal, and the $\beta_{1}$ coefficient is not significantly different from zero. Reassuringly, the number of past campaigns affects a proposal’s success only if the NGO campaigns on a topic being discussed at the AGM ($\beta_{2}=0$ but $\beta_{3}>0$): such campaigns increase a proposal’s success by 8% on average.
@@ 7 Discussion
Our results demonstrate that NGOs strategically use the timing of their campaigns to influence firms’ decisions. In particular, as media coverage is higher during AGMs, NGOs are likely to wait for the AGM date to launch their campaigns against target firms’ practices or strategies (Result 1). By doing so, NGOs prompt reactions from key stakeholders, such as shareholders, who are more likely to file a related proposal at the following AGM, and consumers, as the likelihood of boycotts increases if NGOs campaign on the AGM date (Result 2). The goal of these campaigns is to alter a firm’s strategy by damaging its reputation and showing stakeholders’ disappointment (Bakos and Dellarocas, 2011; Kitzmueller and Shimshack, 2012).
@@
Therefore, NGOs face a trade-off, which they navigate by adapting their strategies over time. NGOs with limited activism histories tend to campaign on the AGM date, while more established NGOs prefer targeting firms just before the AGM to directly influence shareholder votes (Result 4). This pattern is consistent with the idea of using campaigning time as a signaling device. One interpretation of this result is that, as NGOs gain more experience with campaigning—and potentially run multiple campaigns in a year—they accumulate exposure through standard news media and may already have a large social media following. As a result, there may be less pressure to campaign specifically on the AGM date. Another interpretation is that, as NGOs become more complex organizations over time, media exposure becomes less vital for influencing decision-making at target firms. Disentangling these interpretations suggest several potential avenues for future research, contributing to the broader conversation on civil society lobbying and activism (Bertrand et al., 2021, 2023).
@@
The public may also expect mature NGOs not only to generate negative publicity but also to produce immediate results. Thus, there may be diminishing returns to reputation from consistently targeting AGMs. We show that this practice enables NGOs with a history of targeting AGMs to help related shareholder proposals pass. However, an excessively long history of targeting AGMs, which according to our estimates happens after 13 campaigns on average, decreases support for shareholder proposals (Result 5).
@@
In conclusion, NGOs play a crucial role in increasing information and aligning corporate interests with societal needs, and their growth in recent years is well-documented (e.g., Egorov and Harstad, 2017; Daubanes and Rochet, 2019). However, our understanding of the incentives they face, their relationships with firms and stakeholders, and the strategies available to them remains limited due to the scant data available on NGOs and models to interpret their actions. This paper offers a detailed analysis of NGO campaigns to uncover these trade-offs and incentives by leveraging the exogeneity of the corporate calendar to NGO’s goals. Similar approaches have been successfully used to identify the preferences of other stakeholders and may help reframe the role of AGMs not just as voting venues but also as forums where firms engage with and negotiate alongside outside stakeholders in ways that are not often observed by researchers but can be inferred from the underlying system of incentives (Fioretti et al., 2023). We also acknowledge that NGOs not only campaign but also collaborate with corporations (Dahan et al., 2010). Future research could capitalize on similar exogenous variation to that employed in this paper to quantitatively study the incentives for collaboration.
@@
By focusing on the timing of NGO campaigns, we demonstrate how this seemingly innocuous decision has significant implications for both targeted firms and the success of future initiatives. We show that, similar to for-profit organizations, NGOs are influenced by their reputation, which affects the timing of the information presented to stakeholders and, consequently, their responses. Thus, systematically studying NGOs’ objectives and how they evolve over time is essential for understanding whether and how their activism can help firms internalize firms’ externalities.
@@ 8 Conclusion
This paper examines the timing and motives behind NGO campaigns, revealing that these campaigns are not randomly timed but often strategically target firms’ Annual General Meetings. While this timing does not affect votes on environmental and social proposals, which are cast beforehand, it enhances the NGO’s visibility, leading to greater consumer reactions and increasing the likelihood of related shareholder proposals in the following year. Mature NGOs utilize the reputation similarly built over the years to influence current firm decisions by supporting relevant shareholder proposals. Thus, this paper contributes by empirically showing that an NGO’s objectives vary throughout its life, indicating that NGOs pursue different outcomes for targeted firms, their stakeholders, and society at large as they gain visibility, thereby fostering diverse discussions within companies and, potentially, in the broader public arena.
@@ Appendix A Omitted Tables
Table A2: Campaign timing and characteristics of target firms
Characteristics of Target Firm
@@ Appendix B Omitted Figures
Figure B6: NGO experience and the timing of NGO campaigns – Alternative set of Fixed Effects
Note: This figure plots the estimated coefficients from the regressions of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The regression includes replaces NGO-by-year fixed effects with separate fixed effects for each NGO and each year. Hence the coefficients $\{\hat{\beta}_{\tau}\}_{\tau}$ in (5) is also identified from variation in NGO experience over years. Standard errors are clustered at the NGO level.
@@
Note: This figure plots the estimated coefficients from the regressions (5) of the distance between NGO campaigns and the targeted firm’s closest AGM (in months) on dummies indicating the experience of the NGO, measured by the number of campaigns launched before. The panel focuses on NGOs whose first campaign was within the two weeks following the AGM. Standard errors are clustered at the NGO level.
@@
Note: This figure plots the estimated coefficients from regressing the target firm’s market cap (in deciles to control for the fact that the tails of the market cap distribution have thickened in recent years) on dummies indicating the experience of the NGO, measured by the number of campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 1,338.
@@
Note: This figure plots the estimated coefficients from regressing either a target firm’s environmental score (top panel) or its social score (bottom panel) on dummies indicating the experience of the NGO, measured by either the number of environmental (top panel) or social (bottom panel) campaigns it had launched in previous year as in (5). Standard errors are clustered at the firm level. The number of observations is 636 and 698, respectively.